OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (5, -5) with 0° ≤ θ < 360°.

OpenStudy (nikvist):

\[x=5,y=-5\]\[r=\sqrt{x^2+y^2}=5\sqrt{2}\]\[\cos{\theta}=x/r=1/\sqrt{2}\quad,\quad\sin{\theta}=x/r=-1/\sqrt{2}\quad\Rightarrow\quad\theta=7\pi/4=315^{\circ}\]

OpenStudy (anonymous):

Thank you!